OpenStudy (anonymous):

A right cylinder has a height of 5 inches and a radius of 7 inches. What is its volume in terms of PI ? A. 35 cubic inches B. 175 cubic inches C. 245 cubic inches D. 12 cubic inches E. 875 cubic inches

Directrix (directrix):

C. 245 π cubic inches Volume = Bh where B is the area of the base and h is the height of the cylinder V = pi * r^2*h V = pi * 7^2 * 5 V = 245 pi cubic inches